7. If an economy is consuming $19.5 trillion out of its $21.5 trillion, then what is this economy’s MPS?
Answer #1

MArginal Propensity to Consume (MPC) = Change in Consumption/Change in income = 19.5/21.5 = 0.91

Also MPC+MPC = 1

MPS = 1-MPC = 1-0.91 = 0.09